### v1

@@ 1 Introduction
Table 1: Summary of Prior Works on Ephemeris Design for Lunar Navigation Satellites (OE: Osculating Elements)
@@
[7] Chebyshev ELFO (12 hr), LLO
[11] OE + Empirical Acceleration, Chebyshev ELFO (23.95 hr)
[12] OE + Chebyshev ELFO (24.37 hr)
[9] OE + Chebyshev ELFO (12, 18, 24 hr), DRO
This work OE + Chebyshev + Fourier OE (Polynomial + Fourier) ELFO (12, 24, 30 hr), Polar
@@
This paper introduces a unified framework for ephemeris and almanac design tailored for LunaNet satellites. The proposed ephemeris representation integrates osculating orbital elements with Chebyshev polynomials and Fourier series to capture both short- and mid-term orbital perturbations, improving fitting accuracy over longer intervals compared to prior methods. For the almanac, a compact polynomial–Fourier model is developed to represent the long-term evolution of osculating elements and argument of latitude, achieving sub-degree median latitude and longitude errors over 15-day arcs—sufficient for warm-start satellite visibility identification. We validate the ephemeris and almanac generation across multiple orbit regimes: 12-, 24-, and 30-hour ELFOs and low-altitude polar orbits.
@@ 2.1 Coordinate Frames and Dynamics
2.2 Orbit Propagation Analysis
@@ 3.1 Ephemeris Representation
where $c_{0}$ is the average value of the function over one period, which will be shared with the Chebyshev polynomials, $u(t)$ is the argument of latitude at time $t$, and $C_{cn}$ and $C_{sn}$ are the Fourier coefficients for the cosine and sine terms, respectively. We use the argument of latitude to capture the periodic effects corresponding to the orbital period. The longer periodic effects which has period of the lunar sidereal period can be sufficiently captured by the polynomial terms.
@@
$\nu(t)=2\arctan\!\left(\sqrt{\frac{1+e_{0}}{1-e_{0}}}\tan\frac{E(t)}{2}\right).$
@@
6Adjust the longitude of the ascending node for precession:
@@
$x^{\mathrm{PA}}_{0}(t)$ $=x^{p}(t)\cos\lambda(t)-y^{p}(t)\cos i_{0}\sin\lambda(t)$
$y^{\mathrm{PA}}_{0}(t)$ $=x^{p}(t)\sin\lambda(t)+y^{p}(t)\cos i_{0}\cos\lambda(t)$
$z^{\mathrm{PA}}_{0}(t)$ $=y^{p}(t)\sin i_{0}$
@@
$\dot{x}^{\mathrm{PA}}_{0}(t)$ $=-x^{p}(t)\sin\lambda(t)\dot{\lambda}(t)-y^{p}(t)\cos i_{0}\cos\lambda(t)\dot{\lambda}(t)$
@@
$\dot{y}^{\mathrm{PA}}_{0}(t)$ $=x^{p}(t)\cos\lambda(t)\dot{\lambda}(t)-y^{p}(t)\cos i_{0}\sin\lambda(t)\dot{\lambda}(t)$
@@
$\xi^{\mathrm{PA}}(t)=\xi^{\mathrm{PA}}_{0}(t)+\Delta\xi^{\mathrm{PA}}(t).$
$\dot{\xi}^{\mathrm{PA}}(t)=\dot{\xi}^{\mathrm{PA}}_{0}(t)+\Delta\dot{\xi}^{\mathrm{PA}}(t).$
@@ 3.2 Almanac Representation
where $\xi(t)$ denotes the osculating orbital element at time $t$; $\beta_{k}^{\xi}$ are the polynomial coefficients of order $k$; and $C_{c}^{\xi}$ and $C_{s}^{\xi}$ are the Fourier coefficients for the cosine and sine terms, respectively. The frequency $\omega$ is chosen according to the dominant periodicity observed in the Fourier analysis (Section 2): for the semi-major axis, the orbital period $T_{\mathrm{orb}}$ captures short-period variations, whereas for the remaining elements ($e,i,\Omega,\omega,M,u$), the lunar sidereal period $T_{\mathrm{sid}}=27.321661\penalty 10000\ \text{days}$ captures longer-period modulations driven by third-body and gravitational coupling effects.
@@
for each osculating element and $u$.
@@
$\omega(t)$ $=\beta_{0}^{\omega}+\beta_{1}^{\omega}t+\big[C_{c}^{\omega}\cos(\omega_{sid}t)+C_{s}^{\omega}\sin(\omega_{sid}t)\big]$
@@
$u(t)$ $=\omega(t)+\nu(t)+\beta_{0}^{u}+\beta_{1}^{u}t$
@@ 4.1.1 Ephemeris Fitting
    Using Algorithm 1 (lines 1–6), propagate the nominal two-body motion to obtain the predicted state $\mathbf{X}^{\mathrm{PA}}_{0}(t_{m})$ and argument of latitude $u_{m}$ at each node $t_{m}$.
@@
    $\Delta\mathbf{X}^{\mathrm{PAI}}(t_{m})$ $=\mathbf{X}^{\mathrm{PAI}}(t_{m})-\mathbf{X}^{\mathrm{PA}}_{0}(t_{m})$ (6)
@@
    $\min_{\bm{\alpha}}\sum_{m=1}^{N_{T}}\left\|\Delta\mathbf{X}^{\mathrm{PAI}}(t_{m})-T_{m}\bm{\alpha}\right\|_{2}^{2},$ (8)
    where $T_{m}$ denotes the $m$th row of $T$.
@@ 4.1.2 Almanac Fitting
    $\min_{\bm{\alpha}_{a}}\sum_{m=1}^{N_{T}}\|\Delta a(t_{m})-T_{a,m}\bm{\alpha}_{a}\|_{2}^{2}.$ (12)
@@
    Repeat Steps 3–5 for each remaining element $\xi\in\{e,i,\omega,\Omega\}$ using
    $T_{\xi}$ $=$ (13)
@@
    Evaluate the mean anomaly $M_{0}(t_{m})$ using Algorithm 2 (lines 1–4) and compute $\Delta M=M(t_{m})-M_{0}(t_{m})$. Fit $\bm{\alpha}_{M}$ using the same design matrix $T_{M}=T_{\xi}$.
@@
    Finally, compute the argument of latitude residuals $\Delta u=u(t_{m})-[\omega(t_{m})+\nu(t_{m})]$, and solve for $\bm{\alpha}_{u}$ using $T_{u}=T_{\xi}$.
@@ 5.1 Simulation Setup
30 hr ELFO (LCRNS) 1.12E-04 1.86E-02 654 19 5.02E-03 2.68E-02 893 18 3.17E-01 6.98E-01 893 13
24 hr ELFO (LCNS) 5.37E-04 2.80E-02 816 19 5.53E-02 2.12E-01 893 17 7.63E-01 1.63E++$0$ 882 13
12 hr ELFO (LNSS) 1.30E-02 1.00E-01 882 16 2.44E-01 9.01E-01 893 15 2.28E++$0$ 5.32E++$0$ 888 12
6 hr Polar Orbit 7.41E-05 1.60E-02 579 17 8.32E-05 1.61E-02 837 29 2.06E-01 4.43E-01 889 17
@@ 5.2 Ephemeris Fitting Results for the 30-hour ELFO (LCRNS)
Consistent with prior work, augmenting Chebyshev polynomials with osculating elements yields markedly better accuracy at a fixed bit budget than Chebyshev-only fits. Moreover, for fitting ranges $\geq\!240$ min, adding a short Fourier basis to Chebyshev + elements further reduces both position and velocity errors within the 900-bit constraint. The Fourier terms efficiently capture periodic components of the perturbations, allowing lower Chebyshev orders and smaller coefficient ranges, hence fewer bits for comparable or improved accuracy. For short arcs (120 min or less), the Chebyshev + elements model often provides the best accuracy per bit, as the additional Fourier term contributes limited incremental benefit over short durations.
@@
Fig. 7 and 7 compare bit allocations at 360 min for the Chebyshev + OE + Fourier and Chebyshev + OE representations, respectively. While the Fourier model introduces roughly 120 bits for six Fourier coefficients, it enables a reduction in Chebyshev order and coefficient ranges, yielding a better overall accuracy–bit trade.
@@ 5.3 Ephemeris Fitting Results for Other Orbits
Figure 9: Data size versus 95th-percentile position and velocity errors for the 6-hour polar orbits.
@@ 6.1 Almanac Fitting Results for the 30-hour ELFO (LCRNS)
Fig. 10 shows the least-squares almanac fits over 15 days for the residuals of the osculating elements $\{\Delta a,\Delta e,\Delta i,\Delta\Omega,\Delta\omega,\Delta M,\Delta u\}$. The model captures the dominant secular and periodic trends, but does not fully track the highest-frequency content in the semi-major axis, which manifests as along-track (tangential) phase error—visible in the mean anomaly and argument-of-latitude traces.
@@
Fig. 12 summarizes latitude/longitude errors of the fitted satellite orbits. Median and 99th-percentile errors remain within $1^{\circ}$ and $10^{\circ}$, respectively, over 15 days. During periods visible from the lunar south pole (elevation $5^{\circ}$), the latitude/longitude errors are typically within $3^{\circ}$, which is sufficient for discriminating visible satellites at warm start.
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 12: Latitude and longitude errors for the 30-hour ELFO (LCRNS) almanac over 15 days. Blue: epochs visible from the lunar south pole (elevation $\geq 5^{\circ}$); red: not visible.
@@ 6.2 Almanac Fitting Results for Other Lunar Orbits
Fig. 13 show 15-day latitude/longitude errors for the 6-hour polar orbit. Table 3 summarizes all four cases, including the 3 ELFOs.
For ELFOs, latitude/longitude errors increase as orbital period decreases (smaller semi-major axis), with LNSS showing the largest errors. Higher orbital speeds and stronger short-period content amplify high-frequency element variation, which a low-order almanac cannot fully capture. However, these larger errors arise predominantly near perilune when the satellite is often not visible from the south pole; during visible periods, the 99th-percentile latitude/longitude errors remain within $6.5^{\circ}$ across all ELFOs, which is adequate for acquisition and visible-satellite discrimination.
The polar orbit exhibits smaller and more uniform latitude/longitude errors than ELFOs. Given its near-circular geometry, the below-horizon and above-horizon error distributions are similar.
Figure 13: Latitude/longitude errors for the 6-hour polar-orbit almanac over 15 days.
Table 3: Summary of almanac errors during visibility from the lunar south pole (elevation $\geq 5^{\circ}$).
Orbit Position Error [km] Velocity Error [m/s] Latitude Error [deg] Longitude Error [deg]
@@
30 hr ELFO (LCRNS) 85.9 555.8 738.2 2.45 60.7 126.5 0.46 1.29 1.89 0.57 1.64 1.74
24 hr ELFO (LCNS) 169.3 594.0 783.2 4.69 101.5 161.5 0.56 1.57 2.56 0.90 1.79 2.40
12 hr ELFO (LNSS) 94.3 1006.4 1175.4 17.1 315.1 384.3 0.63 5.11 6.25 1.48 2.92 3.37
6 hr Polar Orbit 5.29 154.8 197.3 41.9 60.8 70.2 0.88 2.28 2.95 0.11 1.90 10.72

### v2

@@ 1 Introduction
Table 1: Summary of Prior Works on Ephemeris Design for Lunar Navigation Satellites (OE: Osculating Elements, EA: Empirical Acceleration, Cheby: Chebyshev)
@@
[2] Cheby ELFO (12 hr), LLO
[1] OE + EA, Cheby ELFO (23.95 hr)
[6] OE + Cheby ELFO (24.37 hr)
[16] OE + Cheby ELFO (12, 18, 24 hr), DRO
This work OE + Cheby + Fourier OE (Polynomial + Fourier) ELFO (12, 24, 30 hr), Polar
@@
This paper introduces a unified framework for ephemeris and almanac design for LunaNet satellites. The proposed ephemeris representation integrates osculating orbital elements with Chebyshev polynomials and Fourier series to capture both short- and mid-term orbital perturbations, improving fitting accuracy over longer intervals compared to prior methods. For the almanac, a compact polynomial–Fourier model is developed to represent the long-term evolution of osculating elements and argument of latitude, achieving sub-degree median latitude and longitude errors over 15-day arcs—sufficient for warm-start satellite visibility identification. We validate the ephemeris and almanac generation across multiple orbit regimes: 12-, 24-, and 30-hour ELFOs and low-altitude polar orbits.
@@ 2.1 Coordinate Frames and Dynamics
2.2 Target Orbits
Four distinct lunar orbit regimes are considered in this study: a 30-hour Elliptical Lunar Frozen Orbit (ELFO), a 24-hour ELFO, a 12-hour ELFO, and a 6-hour near-circular polar orbit. The ELFO configurations correspond to the reference and target orbits for the Lunar Communication Relay and Navigation System (LCRNS) [15], the Lunar Communication and Navigation System (LCNS) [8], and the Lunar Navigation Satellite System (LNSS) [11], respectively. The polar orbit represents a notional design for the LANS interoperability demonstration mission [8].
The initial orbital elements for these orbits are summarized in Table 2. These orbits span a representative range of lunar navigation geometries, covering both elliptical frozen and low-altitude polar regimes, and serve as test cases for the proposed ephemeris and almanac design framework.
Table 2: Initial Conditions for Target Lunar Orbits ($a$: semi-major axis, $e$: eccentricity, $i$: inclination, $\Omega$: longitude of ascending node, $\omega$: argument of periapsis, $\nu$: true anomaly)
Orbit Period Initial Epoch Frame $\bm{a}$ (km) $\bm{e}$ $\bm{i}$ (deg) $\bm{\Omega}$ (deg) $\bm{\omega}$ (deg) $\bm{\nu}$ (deg)
LCRNS (ELFO) 30 hr 2027-03-01 00:00:00 UTC PAI 11315.94 0.692 59.373 321.019 92.494 0.000
LCNS (ELFO) 24 hr 2027-01-01 00:00:00 TDB PAI 9748.14 0.700 50.638 91.603 94.344 2.344
LNSS (ELFO) 12 hr 2027-01-01 00:00:00 TDB PAI 6541.40 0.600 62.940 304.170 90.013 0.000
Polar 6 hr 2027-01-01 00:00:00 TDB PAI 3870.00 0.000 89.986 34.191 67.166 7.026
@@ 3.1 Ephemeris Representation
where $c_{0}$ is the average value of the function over one period, which will be shared with the Chebyshev polynomials, $u(t)$ is the argument of latitude at time $t$, and $C_{cn}$ and $C_{sn}$ are the Fourier coefficients for the cosine and sine terms, respectively. We use the argument of latitude to capture the periodic effects corresponding to the orbital period. The longer periodic effects which has period of the lunar sidereal period can be sufficiently captured by the Chebyshev polynomial terms.
@@
$\nu(t)=2\arctan\!\left(\sqrt{(1+e_{0})/(1-e_{0})}\tan(E(t)/2)\right).$
@@
6Adjust the longitude of the ascending node:
@@
${\color[rgb]{0,0,0}x^{\mathrm{PA}}_{\mathrm{OE}}}(t)$ $=x^{p}(t)\cos\lambda(t)-y^{p}(t)\cos i_{0}\sin\lambda(t)$
${\color[rgb]{0,0,0}y^{\mathrm{PA}}_{\mathrm{OE}}}(t)$ $=x^{p}(t)\sin\lambda(t)+y^{p}(t)\cos i_{0}\cos\lambda(t)$
${\color[rgb]{0,0,0}z^{\mathrm{PA}}_{\mathrm{OE}}}(t)$ $=y^{p}(t)\sin i_{0}$
@@
${\color[rgb]{0,0,0}\dot{x}^{\mathrm{PA}}_{\mathrm{OE}}}(t)$ $=-x^{p}(t)\sin\lambda(t)\dot{\lambda}(t)$
$-y^{p}(t)\cos i_{0}\cos\lambda(t)\dot{\lambda}(t)$
@@
${\color[rgb]{0,0,0}\dot{y}^{\mathrm{PA}}_{\mathrm{OE}}}(t)$ $=x^{p}(t)\cos\lambda(t)\dot{\lambda}(t)$
$-y^{p}(t)\cos i_{0}\sin\lambda(t)\dot{\lambda}(t)$
@@
$\xi^{\mathrm{PA}}(t)={\color[rgb]{0,0,0}\xi^{\mathrm{PA}}_{\mathrm{OE}}}(t)+\Delta\xi^{\mathrm{PA}}(t),\ \dot{\xi}^{\mathrm{PA}}(t)={\color[rgb]{0,0,0}\dot{\xi}^{\mathrm{PA}}_{\mathrm{OE}}}(t)+\Delta\dot{\xi}^{\mathrm{PA}}(t)$
@@ 3.2 Almanac Representation
where $\xi(t)$ denotes the osculating orbital element at time $t$; $\beta_{k}^{\xi}$ are the polynomial coefficients of order $k$; and $C_{c}^{\xi}$ and $C_{s}^{\xi}$ are the Fourier coefficients for the cosine and sine terms, respectively. The frequency $\omega$ is chosen according to the dominant periodicity observed in the Fourier analysis (Section 2): for the semi-major axis, the orbital period $T_{\mathrm{orb}}$ captures short-period variations, whereas for the remaining elements ($e,i,\Omega,M,u$), the lunar sidereal period $T_{\mathrm{sid}}=27.321661\penalty 10000\ \text{days}$ captures longer-period modulations driven by third-body and gravitational coupling effects.
@@
for each osculating element and $u$. Note that the argument of periapsis $\omega$ is not corrected since the almanac parameters of $u=\nu+\omega$ can capture the fitting residuals.
@@
$u(t)$ $={\color[rgb]{0,0,0}\nu(t)+\beta_{0}^{u}+\beta_{1}^{u}t}$
@@ 4.1.1 Ephemeris Fitting
    Using Algorithm 1 (lines 1–6), propagate the nominal two-body motion to obtain the predicted state $\hat{\mathbf{X}}^{\mathrm{PA}}_{\mathrm{OE}}(t_{m})$ and argument of latitude $u_{m}$ at each node $t_{m}$.
@@
    $\Delta\mathbf{X}^{\mathrm{PAI}}(t_{m})$ $=\mathbf{X}^{\mathrm{PAI}}(t_{m})-{\color[rgb]{0,0,0}\hat{\mathbf{X}}^{\mathrm{PA}}_{\mathrm{OE}}(t_{m})}$ (6)
@@
    $\min_{\bm{\alpha}}\sum_{m=1}^{N_{T}}\left\|\Delta\mathbf{X}^{\mathrm{PAI}}(t_{m})-{\color[rgb]{0,0,0}\mathbf{T}_{m}}\bm{\alpha}\right\|_{2}^{2},$ (8)
    where $\mathbf{T}_{m}$ denotes the $m$th row of $\mathbf{T}$.
@@ 4.1.2 Almanac Fitting
    $\min_{\bm{\alpha}_{a}}\sum_{m=1}^{N_{T}}\|\Delta a(t_{m})-{\color[rgb]{0,0,0}\mathbf{T_{a,m}}}\bm{\alpha}_{a}\|_{2}^{2}.$ (12)
    where
@@
    Repeat Step 5 for each remaining element $\xi\in\{e,i,\Omega\}$ using
    $\mathbf{T}_{\mathrm{siderial}}$ $=$ (13)
@@
    Compute the predicted mean anomaly $\hat{M}_{0}(t_{m})$ using Algorithm 2 (lines 1–4) and compute $\Delta M=M(t_{m})-\hat{M}_{0}(t_{m})$. Fit $\bm{\alpha}_{M}$ using design matrix $\mathbf{T}_{\mathrm{sidereal}}$.
@@
    Compute the predicted true anomaly $\hat{\nu}(t_{m})$ using fitted eccentricity $\hat{e}(t_{m})=(\mathbf{T}_{sideral})_{m}\bm{\alpha}_{e}$ and mean anomaly $\hat{M}(t_{m})=M_{0}+(\mathbf{T}_{sidereal})_{m}\bm{\alpha}_{M}$
  9. 9.
    Finally, compute argument of latitude residuals $\Delta u(t_{m})=\omega(t_{m})+\nu(t_{m})-\hat{\nu}(t_{m})$, and solve for $\bm{\alpha}_{u}$ using $\mathbf{T}_{\mathrm{sidereal}}$.
@@ 5.1 Simulation Setup
30 hr ELFO 1.12E-04 1.86E-02 654 19 5.02E-03 2.68E-02 893 18 3.17E-01 6.98E-01 893 13
24 hr ELFO 5.37E-04 2.80E-02 816 19 5.53E-02 2.12E-01 893 17 7.63E-01 1.63E++$0$ 882 13
12 hr ELFO 1.30E-02 1.00E-01 882 16 2.44E-01 9.01E-01 893 15 2.28E++$0$ 5.32E++$0$ 888 12
6 hr Polar 7.41E-05 1.60E-02 579 17 8.32E-05 1.61E-02 837 29 2.06E-01 4.43E-01 889 17
@@ 5.2 Ephemeris Fitting Results for the 30-hour ELFO (LCRNS)
Consistent with prior work, augmenting Chebyshev polynomials with osculating elements yields markedly better accuracy at a fixed bit budget than Chebyshev-only fits. Moreover, for fitting ranges $\geq\!240$ min, adding a Fourier basis to Chebyshev + elements further reduces both position and velocity errors within the 900-bit constraint. The Fourier terms efficiently capture periodic components of the perturbations, allowing lower Chebyshev orders and smaller coefficient ranges, hence fewer bits for comparable or improved accuracy. For short arcs (120 min or less), the Chebyshev + elements model often provides the best accuracy per bit, as the additional Fourier term contributes limited incremental benefit over short durations.
@@
Fig. 6 and 6 compare bit allocations at 360 min for the Chebyshev + OE + Fourier and Chebyshev + OE representations, respectively. While the Fourier model introduces roughly 120 bits for six Fourier coefficients, it enables a reduction in Chebyshev order and coefficient ranges, yielding a better overall accuracy–bit trade.
@@ 5.3 Ephemeris Fitting Results for Other Orbits
Figure 9: Data size versus 95th-percentile position and velocity errors for the 6-hour polar orbits. Blue: Chebyshev only; green: Chebyshev + osculating elements; red: Chebyshev + osculating elements + Fourier. Points with black outlines satisfy both error targets (10 m, 2.5 mm/s, horizontal dashed line) within 900-bits (vertical dashed line)
@@ 6.1 Almanac Fitting Results for the 30-hour ELFO (LCRNS)
Fig. 10 shows the least-squares almanac fits over 15 days for the residuals of the osculating elements $\{\Delta a,\Delta e,\Delta i,\Delta\Omega,\Delta M,\Delta u\}$. The model captures the dominant secular and periodic trends, but does not fully track the highest-frequency content in the semi-major axis, which manifests as along-track (tangential) phase error—visible in the mean anomaly and argument-of-latitude traces.
@@
Refer to caption Refer to caption Refer to caption
@@
Figure 12: Latitude, longitude, range, and Doppler shift estimation errors with respect to the South Pole user using the fitted almanac for 30-hour ELFO (LCRNS). Only visible points from the South Pole are plotted for longitude, range, and Doppler shift.
Fig. 12 summarizes the estimated elevation, azimuth, range, and Doppler shift of the fitted satellite orbits. The predicted elevation is within 6 degrees of error, and can predict the visibility (elevation $\geq 5$ deg) of the satellite from the South Pole user correctly for 99.4 % of the timesteps (except for 0.6% of the time where the satellites are predicted visible although they are invisible). The predicted azimuth error with respect to the South Pole is within $\pm 2$ degrees for the timesteps when the satellites are visible. Together, the almanac provides good accuracy to judge visibility and approximately locate the satellite direction. The range and Doppler shift estimation error with respect to the South Pole is within 200 km and 104.7 Hz in the 99th percentile, compared to the true range and Doppler shifts that take values between 6733 to 17940 km and -5320 Hz to 5147 Hz, respectively, providing information to limit the search space for acquisition.
@@ 6.2 Almanac Fitting Results for Other Lunar Orbits
Table 4 summarizes the almanac fitting results for all four cases.
Table 4: Summary of almanac error for 15 days. For azimuth, range, and Doppler shift, the errors are the values when satellites are visible from the South Pole user.
Orbit Elevation Error [deg] Azimuth Error (Visible) [deg] Range Error (Visible) [km] Doppler Error (Visible) [Hz]
@@
30 hr ELFO 0.55 3.03 5.07 0.54 1.64 1.73 42.4 165.6 199.2 25.2 83.9 104.7
24 hr ELFO 0.70 3.90 5.98 0.85 1.54 1.88 59.8 187.9 243.2 46.3 136.7 174.8
12 hr ELFO 1.29 10.85 12.65 1.38 2.28 2.54 65.7 197.4 219.2 141.0 322.5 357.4
6 hr Polar 0.81 2.64 3.64 0.20 2.81 16.84 14.9 60.2 79.2 0.56 1.78 2.24
For ELFOs, the angle, range, and Doppler shift estimation errors increase as the orbital period decreases (i.e., with smaller semi-major axis), with LNSS exhibiting the largest errors. Higher orbital velocities and stronger short-period perturbations amplify high-frequency variations in the orbital elements, which cannot be fully captured by a low-order almanac representation.
Nevertheless, the 12-hour ELFO is able to predict satellite visibility from the lunar south pole for $97.6\%$ of the time. In addition, it effectively constrains the acquisition search space by limiting the range and Doppler shift estimation errors to $243.2$ km and $174.8$ Hz at the 99th percentile, respectively. These errors are small relative to the true ranges and Doppler shifts, which span approximately 5182518289798979 km and 4918-4918 Hz to $4848$ Hz.
The errors for the 6-hour polar orbit are shown in Fig. 13. In this case, both range and Doppler shift errors are significantly smaller than those observed for the ELFOs. This improvement is attributed to reduced perturbations from third-body gravity and lunar gravity anomalies, resulting from the smaller semi-major axis and the near-polar ($90^{\circ}$) inclination of the orbit.
Figure 13: Latitude, longitude, range, and Doppler shift estimation errors with respect to the South Pole user using the fitted almanac for 6-hour polar orbit. Only visible points from the South Pole are plotted for longitude, range, and Doppler shift.